OpenStudy (anonymous):

What is the solution of the linear system of equations? y=2x+8 y=x+5 A. (–3, 2) B. (2, –3) C. (2, 3) D. (3, 2)

OpenStudy (anonymous):

It would be A. (-3,2) Plug it in the the equations. So: 2= 2(-3) + 8 and 2= (-3) + 5 2= -6 +8 2=2 2=2